WAEC Mathematics 2014 Question 46

Question 46

Probability

The probability of an event P happening is 1/5 and that of event Q is 1/4. If the events are independent, what is the probability that neither of them happens?

  1. A.4/5
  2. B.3/4
  3. C.3/5
  4. D.1/20
Show the answer and explanation

Answer: C (3/5)

Explanation

P(neither) = (1 − 1/5)(1 − 1/4) = 4/5 × 3/4 = 3/5.